OpenStudy (anonymous):

A soccer team has played 22 games of which it has won 13 and lost 9. If the team wins its remaining z games and finishes the season with a 70% winning percentage, what is the value of z?

OpenStudy (anonymous):

13/22= 59% (13 + x ) / (22 + x) = .70 or 70% 13 + x = .70 ( 22 + x) 13 + x = 15.4 + .70 x .3x = 2.4 x = 8

OpenStudy (anonymous):

my bad habit lol... x is z